Previous

Calendar

June

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

1

JD153

2

JD154

3

JD155 (rain on and off, cool)
Site Visit (JB, NT): Details , Photos (6)

downloaded from cnr, prof, met, nctc, ucb, uc, soil, tcs, and swapped pc cards at the mrs logger

turned on fans in both licor boxes and the quacker box

checked prof flow rate, looked good

see details for snow depths

span @ 850 psi, N2 @ 1625 psi

snapped phenology photos

brought 2 new co2-in-air cylinders to the trailer from Ramaley (CC282352, CC282376)

4

JD156 (rain on and off, cool)
Site Visit (JB): Details

span @ 840 psi, N2 @ 1525 psi

brought 2 empty N2 cylinders back to Ramaley from the trailer

performed some maintenance on the robot, started summer weekly sampling (see details)

post-calibrated airgas cylinder CC18996, calibrated airgas cylinders CC282352, CC282376 (see details)

5

JD157
Spring Snowstorm

6

JD158

7

JD159

8

JD160

9

JD161

get gray jeep back for rest of summer

Dave Bowling (Utah) at the site this week.

10

JD162

11

JD163

12

JD164 (cool, windy)
Site Visit (SpB, JB, DB, NT, MP, CB, SH, SL): Details , Photos (19)

span @ 750 psi, N2 @ 1290 psi

robot, done=2

at ~10 MST, turned on the "slow" ozone system.

moved the forestcam power directly to the back of the UPS

13

JD165
JB Off

14

JD166

15

JD167

16

JD168
Suburban rental starts (Jeep to Alamosa).

17

JD169
for some (unknown) reason the flow rate in the EC/LI-6262 system dropped slightly for a short time in the morning (seems like a wiring or logger problem since T.ec also dropped at the same time!).

18

JD170
Site Visit (SpB, JB, SL, SH): Details , Photos (22)

span @ 725, N2 @ 1100 psi

grabbed ucb, uc, soil, noah, tc and mrs data

replaced the dessicant vials in the 2.5m LI-7500 (sn 75H-0082)

moved the 2.5m PAR sensors so TRAM can use the mini-tower

brought down empty co2 cylinder (CC18996).

robot, done=1

JB returns

19

JD171 (warm)
Site Visit (SpB, JB): Details , Photos (19)

replaced the dessicant vials in the 2.5m LI-7500 (sn 75H-0082)

moved the 2.5m REBS sensor and configured both PAR and REBs sensors in "tree mount"

at ~11 MST, cleaned krypton hygromter

hauled Maggie's ladder to the shed

replaced all prof inlet filters

swapped desiccant at the LI-6251, blow test and zero good

20

JD172
JB Off

21

JD173
Summer solstice (longest day of the year)

22

JD174

23

JD175
SpB Off

24

JD176
JB returns

25

JD177
Site Visit (JB, SO): Details

afternoon thunder and lightning

tram work

26

JD178
Site Visit (JB, SO, MP, NT, CB, SL, SH): Details

more tram work for JB and SO

swapped 15 flasks at the west robot

MP & Co. started testing the ladder method for wand deployment

27

JD179
Site Visit (JB, SO, SO): Details

more tram setup

28

JD180

29

JD181

30

JD182

1

JD183

2

JD184
SpB Returns

3

JD185
Site Visit (SpB, JB, JH, NT, SH, SL): Details , Photos (32)

span @ 610 psi, N2 @ 675 psi

from 10:15-10:45 MST, changed domes on 26m REBS Rnet sensor

downloaded from cnr, prof, nctc, met, soil, tcs, ucb and uc loggers

changed internal chemicals in the LI-7500 (sn 75H-0084) on the main tower

swapped 15 flasks at the robot

at ~10 MST, increased flow rate in prof system from "9" to "14"

at ~13:00 MST, prof pump (UN828 KNI) at the base of the main tower was dead, we replaced it with a new pump

Cloudy down on the plains; sunny at the tower

4

JD186
CU Holiday (Independence Day)

5

JD187